OpenStudy (anonymous):

Heeelp Me /:

OpenStudy (mew55):

in?

OpenStudy (anonymous):

if f(x)= 1/x & g(x)= 2x, Which of the following statement is true? A)(fog)(1)=2 B)(fog)(0)=3 C)(fog)(2)=1/4 D)(gof)(-2)=7

OpenStudy (anonymous):